how do you do the steps to this problem? 3x=x/2+10

Answer by xcentaur(357)

3x=x/2+10
multiply everything by 2
2(3x)=2(x/2)+2(10)
6x=x+20
6x-x=20
5x=20
x=4


-Prabhat